Reference equation: Real GDP per capita growth rate = Nominal GDP per capita growth rate-inflation rate-Population growth rate This equation is an approximation of the exact rate of growth of GDP per capita, and so it results in some errors when caloulating this rate. However, the smplified equation is both easy to use and results in small error terms when inflation, nominal GDP growth, and population growth are low, and so it is a useful approximation. The table below lists a fictional countrys nominal GOP, real GDP. GDP delator, and population over two years Nominal GOP GDP deflator Real GDP 13 ,050,000 102 1029412 1,005 a. Verity that the real GDP values in the above table are accurate. Instructions: Round your answers to one decimal place b. This countrys real GOP per capita for 2012 is This countrys real GOP per capita for 2013 is С.The growth rate in this countrys real GDP per capita between 2012 and 2013 d. The growth rate in nominal GOP between 2012 and 2013 i The growh rate in the GOP defliator between 2012 and 2013 The growh rate in the population between 2012 and 2013
Answer #1

a. Real GDP in 2012 = (1,000,000 * 100) * 100 = 1,000,000

    Real GDP in 2013 = (1,050,000 * 100) * 102 = 1,029,412

b. Real GDP per capita in 2012 = 1,000,000 / 1,000 = $1,000

    Real GDP per capita in 2013 = 1,029,412 / 1,005 = $1,024.3

c. Growth rate in real GDP per capita between 2012 & 2013 = [(1024.3 - 1000) / 1000] * 100 = 2.43%

d. Growth rate in nominal GDP between 2012 & 2013 = [(1,050,000 - 1,000,000) / 1,000,000] * 100 = 5%

     Growth rate in the GDP deflator between 2012 & 2013 = [(102 - 100) / 100] * 100 = 2%

     Growth rate population between 2012 & 2013 = [(1,005 - 1,000) / 1,000] * 100 = 0.5%